OpenStudy (anonymous):

whats the answer to the following cuz i have tried.... "Find the product and simplify. Assume the variable represents a nonnegative real number" 2times square root of 6x times 3times square root of 8x^8

myininaya (myininaya):

\[2 \cdot \sqrt{6x} \cdot 3 \cdot \sqrt{8x^8}\] ?